# Efficient Generative Model Training via Embedded Representation Warmup

### v2

@@ Appendix A Loss Function Decomposition Analysis
$\mathcal{L}_{\text{diff}}=\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|F_{\bm{\theta}}(\mathbf{z}_{t},t)-\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t})\|^{2}\right],$ (17)
where $F_{\bm{\theta}}(\mathbf{z}_{t},t)$ parameterizes the estimated score function, and $\lambda(t)$ is a time-dependent weighting function.
Our analysis hinges on the idealized existence of a sufficient representation $\mathbf{r}_{t}=R(\mathbf{z}_{t})$ that could statistically decouple the latent space geometry from generation dynamics. Formally, we assume:
@@
Assumption 1 (Sufficient Representation) .
The representation $\mathbf{r}_{t}=R(\mathbf{z}_{t})$ is assumed to be a sufficient statistic for $\mathbf{z}_{t}$ with respect to the data distribution $p(\mathbf{z}_{t})$. This implies that $\mathbf{r}_{t}$ would capture all necessary information about $\mathbf{z}_{t}$ for generation. This strong assumption leads to the hypothetical decomposition of the Fisher score:
$\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t})=\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})+\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}_{t}),$ (18)
where the generation score $\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})$ and the representation score $\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}_{t})$ would be orthogonal under the Fisher-Rao metric:
$\mathbb{E}_{\mathbf{z}_{t}\mid\mathbf{r}_{t}}\left[\langle\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t}),\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}_{t})\rangle\right]=0.$ (19)
This idealized orthogonality would reflect the statistical independence between the generation mechanism and the representation learning process, ensuring that the two components could be optimized independently. However, in a practical neural network, these components are inevitably coupled.
Under this idealized assumption, the ground truth score decomposes into generation and representation components:
@@
$\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t})=\underbrace{\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})}_{\text{Generation Score}}+\underbrace{\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}_{t})}_{\text{Representation Score}}.$ (20)
Parametrizing the model’s score function as $F_{\bm{\theta}}(\mathbf{z}_{t},t)=\mathcal{G}_{\bm{\theta}}(\mathbf{r}_{t})+\mathcal{R}_{\bm{\theta}}(\mathbf{z}_{t})$, where $\mathcal{G}_{\bm{\theta}}$ and $\mathcal{R}_{\bm{\theta}}$ are intended to independently model generation and representation scores, we expand the loss function:
$\mathcal{L}_{\text{diff}}$ $=\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|(\mathcal{G}_{\bm{\theta}}(\mathbf{r}_{t})+\mathcal{R}_{\bm{\theta}}(\mathbf{z}_{t}))-(\nabla\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})+\nabla\log p(\mathbf{r}_{t}))\|^{2}\right]$ (21)
@@
$=\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|\mathcal{R}_{\bm{\theta}}(\mathbf{z}_{t})-\nabla\log p(\mathbf{r}_{t})\|^{2}\right]$
$\quad+\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|\mathcal{G}_{\bm{\theta}}(\mathbf{r}_{t})-\nabla\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})\|^{2}\right]$
$\quad+2\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\langle\mathcal{R}_{\bm{\theta}}(\mathbf{z}_{t})-\nabla\log p(\mathbf{r}_{t}),\mathcal{G}_{\bm{\theta}}(\mathbf{r}_{t})-\nabla\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})\rangle\right].$ (22)
The cross-term would vanish due to the idealized orthogonality condition in Assumption 1, as:
$\mathbb{E}_{\mathbf{z}_{t}\mid\mathbf{r}_{t}}\left[\langle\nabla\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t}),\nabla\log p(\mathbf{r}_{t})\rangle\right]=0.$ (23)
This yields the following decoupled learning objectives, which serve as a conceptual guide for our method:
@@
$\mathcal{L}_{\text{diff}}\approx\underbrace{\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|\mathcal{R}_{\bm{\theta}}(\mathbf{z}_{t})-\nabla\log p(\mathbf{r}_{t})\|^{2}\right]}_{\mathcal{L}_{\text{L2R}}}+\underbrace{\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|\mathcal{G}_{\bm{\theta}}(\mathbf{r}_{t})-\nabla\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})\|^{2}\right]}_{\mathcal{L}_{\text{R2G}}}.$ (24)
Through the above conceptual decomposition, we can approximately split the optimization objective of the diffusion model into two largely independent parts: Representation Learning $\mathcal{R}_{\bm{\theta}}(\mathbf{z}_{t})$ and Generation Inversion $\mathcal{G}_{\bm{\theta}}(\mathbf{r}_{t})$. This provides the intuition for our two-phase training approach, even though perfect orthogonality is not achievable in practice.
Appendix B Analysis Details
B.1 Evaluation details
@@
B.2 Description of pretrained visual encoders
@@
Appendix C Hyperparameter and More Implementation Details
@@
Input dim. 32$\times$32$\times$4 32$\times$32$\times$4 32$\times$32$\times$4
@@
Further implementation details. We implement our model based on the original SiT implementation (Ma et al., 2024). Throughout the experiments, we use the exact same structure as DiT (Peebles & Xie, 2023) and SiT (Ma et al., 2024). We use AdamW (Kingma, 2015; Loshchilov, 2017) with constant learning rate of 1e-4, $(\beta_{1},\beta_{2})=(0.9,0.999)$ without weight decay. To speed up training, we use mixed-precision (fp16) with a gradient clipping. We also pre-compute compressed latent vectors from raw pixels via stable diffusion VAE (Rombach et al., 2022) and use these latent vectors. Because of this, we do not apply any data augmentation, but we find this does not lead to a big difference, as similarly observed in EDM2 (Karras et al., 2024). We also use stabilityai/sd-vae-ft-ema decoder for decoding latent vectors to images. For MLP used for a projection, we use three-layer MLP with SiLU activations (Elfwing et al., 2018). We provide a detailed hyperparameter setup in   Table 6.
@@
Appendix D Evaluation Details
We strictly follow the setup and use the same reference batches of ADM (Dhariwal & Nichol, 2021) for evaluation, following their official implementation. [^2] https://github.com/openai/guided-diffusion/tree/main/evaluations We use 8$\times$×NVIDIA H800 80GB GPUs or for evaluation and enable tf32 precision for faster generation, and we find the performance difference is negligible to the original fp32 precision.

### v3

@@ B.2.2 Marginal and Conditional Distributions
$=\frac{p(\mathbf{z}_{t}\mid\mathbf{z}_{0},t)\,\delta\left(\mathbf{r}-\mathbf{f}_{\text{rep}}\left(\mathcal{D}_{\bm{\theta}_{\text{VAE}}}(\mathbf{z}_{0})\right)\right)\,p(\mathbf{z}_{0})}{p(\mathbf{z}_{t},\mathbf{r},t)}$ (30)
This conditional distribution is the target that our diffusion model seeks to approximate, representing the posterior over clean latents given both the noisy observation and the semantic constraint.
The key insight from equation (27) is that the semantic constraint creates a coupling between $\mathbf{z}_{t}$ and $\mathbf{r}$ through the latent variable $\mathbf{z}_{0}$, despite $\mathbf{z}_{t}$ and $\mathbf{r}$ being conditionally independent given $\mathbf{z}_{0}$. This coupling is what enables semantic-conditioned generation.
The augmented probability space construction embeds the desired semantic knowledge directly into the probabilistic model. The generative task is thus transformed from learning an unconditional reverse process to learning a conditional reverse process, where synthesis is explicitly conditioned on a target semantic concept $\mathbf{r}$. This transformation is fundamental to understanding why our two-phase training approach is theoretically justified.
B.2.3 Semantic Sufficiency and Conditional Independence
The power of the augmented formulation relies on a key assumption about the semantic representation, which we formalize below.
@@ B.2.3 Semantic Sufficiency and Conditional Independence
Assumption 1 (Semantic Sufficiency) .
The semantic representation $\mathbf{r}=\mathbf{f}_{\text{rep}}(\mathbf{x})$ captures sufficient information for the generative task such that, given $\mathbf{r}$, the model possesses all necessary high-level information to synthesize a corresponding sample. Formally, this means that the conditional distribution $p(\mathbf{z}_{0}\mid\mathbf{r})$ concentrates on semantically-consistent latents.
Intuitive Understanding: This assumption embodies the idea that our pretrained representation model $\mathbf{f}_{\text{rep}}$ (e.g., DINOv2) is sufficiently powerful to capture all the high-level, conceptual information needed for generation. To illustrate with an analogy: if $\mathbf{r}$ represents “a golden retriever running on grass,” then semantic sufficiency means that knowing this $\mathbf{r}$ provides the model with all the essential semantic components—the subject (dog), category (golden retriever), action (running), and environment (grass). The model’s remaining task shifts from deciding what to generate to focusing purely on how to generate it: the specific pose, fur details, lighting direction, grass texture, etc.
Latent Space Partitioning: More precisely, we assume there exists a partition of the latent space based on semantic content. The semantic representation $\mathbf{r}$ acts like a clustering label that groups clean latents with identical semantic meaning. We define semantic equivalence classes:
$\mathcal{Z}_{\mathbf{r}}=\{\mathbf{z}_{0}\in\mathcal{Z}:\mathbf{f}_{\text{rep}}(\mathcal{D}(\mathbf{z}_{0}))=\mathbf{r}\}$ (31)
For example, $\mathcal{Z}_{\text{``cat''}}$ might contain latents corresponding to “a crouching Persian cat,” “a rolling orange tabby,” and “a sleeping Siamese cat.” Despite their vastly different visual details, all belong to the same semantic category under $\mathbf{f}_{\text{rep}}$.
Overlap Requirement for Well-Posed Generation: A critical consequence of semantic sufficiency is that for any two clean latents $\mathbf{z}_{0},\mathbf{z}_{0}^{\prime}\in\mathcal{Z}_{\mathbf{r}}$, their respective forward diffusion processes $p(\mathbf{z}_{t}\mid\mathbf{z}_{0},t)$ and $p(\mathbf{z}_{t}\mid\mathbf{z}_{0}^{\prime},t)$ should have significant overlap. This requirement ensures that conditional generation remains well-posed:
  • Without overlap: If semantically similar $\mathbf{z}_{0}$ values produce completely different noisy patterns $\mathbf{z}_{t}$, the model becomes “confused”—it cannot learn a consistent denoising pattern for the semantic class $\mathbf{r}$.
  • With overlap: When $\mathbf{z}_{0}$ values in $\mathcal{Z}_{\mathbf{r}}$ yield similar noisy distributions, the model can learn a unified denoising strategy conditioned on $\mathbf{r}$.
B.2.4 Conditional Score Matching in the Augmented Space
The central idea is to model the score of the joint conditional distribution $p(\mathbf{z}_{0},\mathbf{r}\mid\mathbf{z}_{t},t)$, which naturally decomposes into two meaningful components.
{theoremframe}
Theorem 1 (Decomposition of the Augmented Conditional Score) .
The score of the joint conditional distribution $p(\mathbf{z}_{0},\mathbf{r}\mid\mathbf{z}_{t},t)$ can be decomposed into a sum of two functionally distinct scores:
@@ B.2.4 Conditional Score Matching in the Augmented Space
$\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{0},\mathbf{r}\mid\mathbf{z}_{t},t)=\underbrace{\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{0}\mid\mathbf{z}_{t},\mathbf{r},t)}_{\text{Conditional Generation Score}}+\underbrace{\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}\mid\mathbf{z}_{t},t)}_{\text{Representation Inference Score}}$ (32)
Proof.
We provide a detailed derivation of this fundamental decomposition.
Step 1: Probabilistic factorization. Using the chain rule of conditional probability, we can factorize the joint conditional distribution:
$p(\mathbf{z}_{0},\mathbf{r}\mid\mathbf{z}_{t},t)=p(\mathbf{z}_{0}\mid\mathbf{z}_{t},\mathbf{r},t)\,p(\mathbf{r}\mid\mathbf{z}_{t},t)$ (33)
This factorization is always valid and separates the problem into two components: generating clean latents given both noisy latents and semantic information, and inferring semantic information from noisy latents.
Step 2: Logarithmic transformation. Taking the natural logarithm of both sides of equation (33):
$\log p(\mathbf{z}_{0},\mathbf{r}\mid\mathbf{z}_{t},t)$ $=\log\left[p(\mathbf{z}_{0}\mid\mathbf{z}_{t},\mathbf{r},t)\,p(\mathbf{r}\mid\mathbf{z}_{t},t)\right]$ (34)
$=\log p(\mathbf{z}_{0}\mid\mathbf{z}_{t},\mathbf{r},t)+\log p(\mathbf{r}\mid\mathbf{z}_{t},t)$ (35)
where we used the logarithm property $\log(ab)=\log a+\log b$.
Step 3: Gradient computation. Applying the gradient operator $\nabla_{\mathbf{z}_{t}}$ with respect to the noisy latent $\mathbf{z}_{t}$ to both sides of equation (35):
$\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{0},\mathbf{r}\mid\mathbf{z}_{t},t)$ $=\nabla_{\mathbf{z}_{t}}\left[\log p(\mathbf{z}_{0}\mid\mathbf{z}_{t},\mathbf{r},t)+\log p(\mathbf{r}\mid\mathbf{z}_{t},t)\right]$ (36)
$=\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{0}\mid\mathbf{z}_{t},\mathbf{r},t)+\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}\mid\mathbf{z}_{t},t)$ (37)
where we used the linearity of the gradient operator: $\nabla(f+g)=\nabla f+\nabla g$.
Step 4: Functional interpretation. The resulting decomposition has clear functional meaning:
  • $\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{0}\mid\mathbf{z}_{t},\mathbf{r},t)$ represents the Conditional Generation Score: given both noisy input $\mathbf{z}_{t}$ and semantic target $\mathbf{r}$, how should we move in latent space to increase the likelihood of the clean latent $\mathbf{z}_{0}$?
  • $\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}\mid\mathbf{z}_{t},t)$ represents the Representation Inference Score: given only noisy input $\mathbf{z}_{t}$, how should we move in latent space to increase the likelihood of the semantic representation $\mathbf{r}$?
This completes the proof of the score decomposition in equation (32). ∎
{corollaryframe}
Corollary 1 (Functional Interpretation of Score Components) .
Thm. 1 provides the central theoretical insight of our work. The total learning objective is a linear superposition of two functionally distinct tasks:
  1. 1.
    Conditional Generation Score: The term $\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{0}\mid\mathbf{z}_{t},\mathbf{r},t)$ corresponds to the R2G (Representation-to-Generation) circuit. It addresses the pure synthesis problem: given a noisy latent $\mathbf{z}_{t}$ and the ground-truth semantic concept $\mathbf{r}$, compute the score vector towards the clean latent $\mathbf{z}_{0}$.
  2. 2.
    Representation Inference Score: The term $\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}\mid\mathbf{z}_{t},t)$ corresponds to the L2R (Latent-to-Representation) circuit. It addresses the semantic inference problem: given only a noisy latent $\mathbf{z}_{t}$, compute the score vector that increases the likelihood of the underlying semantic representation being $\mathbf{r}$.
B.2.5 Emergence of the Two-Phase Training Framework from the Theory
A naive attempt to train a single, monolithic network $\mathbf{F}_{\bm{\theta}}$ to approximate the joint score in (32) would re-entangle the two objectives, leading to optimization conflicts. A more principled approach, suggested by the decomposition itself, is a curriculum learning strategy that addresses the two scores in a structured sequence. This naturally gives rise to the ERW framework.
{lemmaframe}
Lemma 1 (Phase 1: Representation Warmup as Boundary Condition Matching) .
The first phase of ERW, the representation warmup, can be interpreted as learning the Representation Inference Score at the clean boundary condition, i.e., at $t=0$.
Proof.
We provide a detailed derivation showing how the warmup phase corresponds to boundary condition matching.
Step 1: Analysis at the boundary condition. At $t=0$, the forward diffusion process gives us $\mathbf{z}_{t}=\mathbf{z}_{0}$ (no noise added). The Representation Inference Score becomes:
$\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}\mid\mathbf{z}_{t},t)\big|_{t=0}=\nabla_{\mathbf{z}_{0}}\log p(\mathbf{r}\mid\mathbf{z}_{0})$ (38)
Step 2: Simplification using the semantic constraint. From equation (19), we have:
$p(\mathbf{r}\mid\mathbf{z}_{0})=\delta\left(\mathbf{r}-\mathbf{f}_{\text{rep}}\left(\mathcal{D}_{\bm{\theta}_{\text{VAE}}}(\mathbf{z}_{0})\right)\right)$ (39)
This is a Dirac delta function, which means the score is not well-defined in the classical sense. However, we can interpret this in terms of the desired functional behavior.
Step 3: Functional interpretation and approximation. In practice, we approximate the deterministic relationship through a learned mapping. The warmup objective is:
$\mathcal{L}_{\mathrm{warmup}}=\mathbb{E}_{\mathbf{x}\sim p_{\text{data}}}\Big[\,\ell_{\mathrm{NT\text{-}Xent}}\big(\mathcal{T}_{\bm{\theta}}(\mathcal{R}_{\bm{\theta}_{\text{L2R}}}(\mathcal{H}_{\bm{\theta}_{\text{VAE}}}(\mathbf{x}))),\;\mathbf{f}_{\text{rep}}(\mathbf{x})\big)\Big]$ (40)
where $\mathcal{R}_{\bm{\theta}_{\text{L2R}}}$ is the L2R circuit that we train to approximate the mapping $\mathbf{z}_{0}\mapsto\mathbf{r}$.
Step 4: Connection to boundary condition. Optimizing NT-Xent at t=0t{=}0 serves as boundary-condition matching for representation alignment. Specifically, we want:
$\mathcal{R}_{\bm{\theta}_{\text{L2R}}}(\mathbf{z}_{0})\approx\mathbf{f}_{\text{rep}}(\mathcal{D}_{\bm{\theta}_{\text{VAE}}}(\mathbf{z}_{0}))=\mathbf{r}$ (41)
This provides a strong "semantic anchor" for the model at $t=0$, ensuring that the L2R circuit learns to extract meaningful semantic representations from clean latents under the same contrastive objective used in Phase 2.
Step 5: Extension to $t>0$. Once the boundary condition is satisfied, the L2R circuit can be expected to generalize to noisy inputs $\mathbf{z}_{t}$ for $t>0$, providing a foundation for the representation inference score at all time steps. ∎
{lemmaframe}
Lemma 2 (Phase 2: Guided Synthesis as Joint Score Optimization) .
The second phase of ERW, guided synthesis, corresponds to learning the full joint score, where the two components from Thm. 1 are learned concurrently under a curriculum.
Proof.
We demonstrate how Phase 2 implements joint score optimization through a carefully designed curriculum.
Step 1: Phase 2 objective decomposition. After the warmup phase, the L2R circuit $\mathcal{R}_{\bm{\theta}_{\text{L2R}}}$ is a competent representation extractor. The Phase 2 total loss is:
$\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{diffusion}}+\lambda_{\mathrm{train}}(k)\cdot\mathcal{L}_{\mathrm{align}}$ (42)
where:
@@ B.2.5 Emergence of the Two-Phase Training Framework from the Theory
$\mathcal{L}_{\mathrm{diffusion}}$ $=\mathbb{E}_{t,\mathbf{z}_{t},\mathbf{z}_{0}}\left[w(t)\left\|\mathbf{F}_{\bm{\theta}}(\mathbf{z}_{t},t)-\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{0}\mid\mathbf{z}_{t},t)\right\|^{2}\right]$ (43)
@@
$\mathcal{L}_{\mathrm{align}}$ $=\mathbb{E}_{\mathbf{z}_{t},\mathbf{r}}\left[\ell_{\mathrm{align}}\bigl(\mathcal{R}_{\bm{\theta}_{\text{L2R}}}(\mathbf{z}_{t},t),\;\mathbf{r}\bigr)\right]$ (44)
Step 2: Connection to the score decomposition. From Thm. 1, the joint conditional score decomposes as:
$\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{0},\mathbf{r}\mid\mathbf{z}_{t},t)=\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{0}\mid\mathbf{z}_{t},\mathbf{r},t)+\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}\mid\mathbf{z}_{t},t)$ (45)
Our Phase 2 objective should be interpreted as shaping these two functional components via practical surrogate losses: the standard diffusion loss $\mathcal{L}_{\mathrm{diffusion}}$ for generation and the alignment loss $\mathcal{L}_{\mathrm{align}}$ for representation, rather than claiming exact equality to the joint score at all times. In practice, we instantiate $\ell_{\mathrm{align}}$ as a contrastive objective (e.g., NT-Xent) with in-batch negatives.
Step 3: Curriculum learning analysis. The training-schedule-dependent weighting $\lambda_{\mathrm{train}}(k)$ creates a curriculum that balances the two objectives:
  • Early in Phase 2 (large $\lambda_{\mathrm{train}}(k)$):
    $\mathcal{L}_{\mathrm{total}}\approx\lambda_{\mathrm{train}}(k)\cdot\mathcal{L}_{\mathrm{align}}+\mathcal{L}_{\mathrm{diffusion}}$ (46)
    The optimization is strongly guided to maintain semantic consistency on noisy inputs, reinforcing the L2R circuit’s ability to extract representations from $\mathbf{z}_{t}$ for $t>0$.
  • Late in Phase 2 (small $\lambda_{\mathrm{train}}(k)$):
    $\mathcal{L}_{\mathrm{total}}\approx\mathcal{L}_{\mathrm{diffusion}}$ (47)
    The L2R circuit is assumed to be robust, and optimization focus shifts to perfecting the full score matching. This allows the R2G circuit to learn the Conditional Generation Score while relying on stable, high-quality representations from the L2R circuit.
Appendix C Analysis Details
C.1 CKNNA Metric Details
@@
C.2 Description of pretrained visual encoders
@@
Appendix D Hyperparameter and More Implementation Details
D.1 Hyperparameter Tuning
  1. (a)
  2. (b)
    To evaluate performance quickly, we reduce the sampling steps from the usual $250$ to $50$ and generate only $10\mathrm{k}$ samples (instead of $50\mathrm{k}$) to compute a preliminary FID score.
@@ D.1 Hyperparameter Tuning
Input dim. $\times32$$\times$4 $\times32$$\times$4 $\times32$$\times$4
@@
Further implementation details. We implement our model based on the original SiT implementation (Ma et al., 2024). Throughout the experiments, we use the exact same structure as DiT (Peebles & Xie, 2023) and SiT (Ma et al., 2024). We use AdamW (Kingma, 2015; Loshchilov, 2017) with constant learning rate of 1e-4, $(\beta_{1},\beta_{2})=(0.9,0.999)$ without weight decay. To speed up training, we use mixed-precision (fp16) with gradient clipping at norm 1.0. We also pre-compute compressed latent vectors from raw pixels via stable diffusion VAE (Rombach et al., 2022) and use these latent vectors. Because of this, we do not apply any data augmentation, but we find this does not lead to a big difference, as similarly observed in EDM2 (Karras et al., 2024). We also use stabilityai/sd-vae-ft-ema decoder for decoding latent vectors to images. For MLP used for a projection, we use three-layer MLP with SiLU activations (Elfwing et al., 2018). We provide a detailed hyperparameter setup in   Table 6.
@@
Appendix E Evaluation Details
[…]